OpenStudy (anonymous):

the angle from a point on the ground to the top of a 168 ft tower is 35 degrees

OpenStudy (anonymous):

a) sin 35 degrees =168/a b)tan 55 degrees= 168/a c) cos 35 =168/a d) cos 55 =a/168

OpenStudy (anonymous):

#3

OpenStudy (anonymous):

@tomo can u help me on this one plzz !

OpenStudy (anonymous):